OpenStudy (anonymous):

A company will need $30,000 in 5 years. to meet this goal they will need to deposit money in an account that pays 7% annual interest compound quarterly. find the amount that will be invested to total $30,000 in 5 years.

OpenStudy (jim766):

A = P ( 1 + r/n)^t*n p is principal, the amount to invest r is rate as a decimal n is number times compounded in a yr t is number of years \[30,000=A(1 + \frac{ .07 }{ 4 })^{5*4}\]

OpenStudy (jim766):

solve for A

OpenStudy (anonymous):

same as: http://openstudy.com/study#/updates/5185addee4b0bc9e008a6df6